mnkjhugy:

Paolo is putting a fence around his pool. The length of the pool is 20 feet, the width is (20 minus x) feet, and the area of the pool is 300 square feet. Recall that area is length multiplied by width. The equation below models the area of the pool.

jhonyy9:

let the length be l and the width be w given l = 20 ft and w = 20-x area = l*w so 300 = l*w what is the question here ?